# Augmenting LLMs to Repair Obsolete Test Cases with Static Collector and Neural Reranker

### v1

@@ V-A2 Human Evaluation on Effectiveness
Since textual metrics focus on measuring the similarity of tokens and AST structure for the given codes, they can not well represent the correctness of repair. To bridge this gap, in this part, we replace the obsolete test cases with the generated repaired ones and manually check whether the replaced test is correct or not, where a correct repair should pass the compilation and keep the intent of the test unchanged.
@@
| Synbciatr | 96.2% | 75.0% (102) | 90.4% (123) |
The column CPR in Tab. II represents the success rate of compilation after repairing the original test case in the repository. As a result, Synbciatr fixes the most compilation errors caused by SynBC. In terms of the two metrics of repairability, Synbciatr both outperforms the baselines. As shown in the column Intent Match, we can observe that Synbciatr correctly repair 75.0% (102/136) and 90.4% (123/136) test cases in alignment with the intent of the ground truth and the original test respectively.
As for specifically, Synbciatr achieves improvements of 248.6% and 9.8% when compared to Ceprot and naiveLLM respectively.
@@ V-B RQ2: Effectiveness of TROCtx
In this research question, we focus on comparing our proposed Synbciatr with Naivellm to evaluate the effectiveness of TROCtx. As introduced in Section I, TROCtx serves as contextual information to reduce hallucinations during the generation of LLM (LLM below all refers to GPT-4).
@@
Common hallucinations. Common hallucinations are caused by the lack of direct contextual information from the change of focal method signature. For the case in Fig. 1, a new class (MountPOptions) is set as the third parameter type. Therefore, LLM falls into common hallucination without the required class contexts of MountPOptions.
Outdated hallucinations. Outdated hallucinations are caused by the lack of implicit contextual information from the focal method body. For the example in Fig. 5, the method findUnknown used in the test is refactored to be accessed from an instance of class BitstreamFormat. Without this specific knowledge, the test generated by LLM still uses the method in an outdated way, which causes hallucinations.
To assess the effectiveness of TROCtx constructed by Synbciatr, we collect all the hallucinations for Naivellm and Synbciatr. As shown in Tab. III, hallucinations occur in both approaches. For the 6 common hallucinations that Naivellm fails, Synbciatr fixes all of them and correctly generates repaired tests with related contexts, indicating that Synbciatr has the general capability to precisely and effectively collect required contexts from the syntactic change of the focal method. Besides, Synbciatr also reduces outdated hallucinations by 25%, in which two cases are fixed with the constructed TROCtx (UsageCtx and EnvCtx specifically).
@@ V-C RQ3: Failure Analysis
Synbciatr is designed to repair the original test case with its intent unchanged. Considering that the ground truth may contain semantic changes such as adding new assertions and altering input values, we further investigate the 13 cases that fail in terms of according to Tab. II. Finally, we summarize three reasons for the failure.
Uses Unimported Classes. For four cases, LLM directly uses classes that are not imported in the file of the test method when generating codes. The key reason is that we focus on collecting contexts by analyzing the syntactic change of the focal method at the method level. Without analyzing the focal class and changes in the methods, the constructed contexts by Synbciatr may be inadequate and result in failure. Fortunately, this type of failure is easy to fix.
Complex Focal Changes. For eight cases, LLM fails to correctly generate the expected test with limited contexts as the changes of focal methods in class-level or repo-level upgrades are complex. Specifically, the complex focal change leads to outdated hallucinations and incorrect invocations. This type is challenging to resolve because of the difficulty of collecting implicit contexts and the limited capability of current LLMs.
Fails to construct TROCtx. Specifically for one case, We find that Language Server fails to provide intelligent features for repositories containing configuration error, which results in the failure of Synbciatr to construct TROCtx. Overall, it is hard even for developers to manually repair tests in these repositories, in which they can not infer related contexts in IDEs either.
@@ V-D RQ4: Efficiency
Compared to Naivellm, Synbciatr adds the steps to construct TROCtx. When querying LLM to repair the test, the prompt of Synbciatr is longer with the constructed TROCtx. In this research question, we evaluate the time efficiency and the token count of Synbciatr compared to Naivellm.
In particular, we divide the process of Synbciatr into constructing TROCtx and querying LLM for repair. As shown in Fig. 6, the time cost of Synbciatr mainly depends on constructing TROCtx, in which the response time of LLM API of two approaches is close. We notice that the Language server may fail to synchronize old repositories due to connection timeout for missing dependencies, which results in the outliers with much longer time. In our design, Synbciatr trades time for accuracy. Besides, the average time of Synbciatr falls into 10s to 30s, which is acceptable for developers since test repair is not an interactive task.
In terms of the token cost of Synbciatr, we observe that the tokens of the prompt roughly doubled in number after being augmented with TROCtx from Fig. 7. This evidence also indicates that the TROCtx constructed by Synbciatr is precise.
VI Threats to Validity
In this section, we discuss the possible threats to validity from external and internal aspects respectively.
External Validity. The main threats to external validity come from the evaluation dataset. We reuse the existing dataset, which may not be representative of all possible real-world syntactic changes, so we check and augment the dataset to collect diverse samples from related commits. Besides, it is inevitable to avoid data leakage from popular open-source projects in GitHub. These projects widely use standard design patterns [36], which can be learned by LLM during training. Although experimental results show that Naivellm correctly repairs some of the tests with data leakage, it does not affect the effectiveness evaluation of Synbciatr. According to the analysis in RQ2, Synbciatr outperforms Naivellm in cases where Naivellm fails for hallucinations.
@@ VII-A Production-test Co-evolution
There are two types of obsolete test cases that need repair, one is the GUI-oriented event sequence test case, the other is the code-oriented method test case. Nowadays, many researchers studied the automated repair of GUI test cases, especially on Android applications [38], [11]. However, for code-oriented method test cases, there are limited studies focused on repairing code-oriented method test cases automatically. The biggest difference between them is that the searching space for the GUI-oriented test event is more oblivious, which can be obtained by traversing the related GUI widget trees, while the code-oriented method test case has a larger searching space in the whole repository.
For code-oriented method test case repairing, Hu et al. [12] proposed the first transformer-based approach to update obsolete tests with two stages, identifying and updating, which is the SOTA work in this area. However, it is based on pre-trained models with fewer parameters and lacks contextual information. Compared to this, our work uses larger language models (LLMs) with automatically constructed contexts from the repository.
@@ VII-B LLM-based Code Generation
Automated code generation with LLM brings huge improvements in production efficiency. Repo-level code generation represents the task of generating codes based on a broader context of the repository [39]. It is challenging due to the lack of domain-specific knowledge, which results in hallucination [40], [41]. Several studies leverage RAG to improve the performance of LLM in specific code tasks by providing similar codes or results into the query prompt [39], [25]. However, the contexts are retrieved based on simple metrics such as textual similarity without code semantics. Recently, some works have focused on improving the capability of LLM by static analysis [42], [18]. Specifically, monitor-guided decoding (MGD) [18] is a novel approach to bring IDE-assistance from developers to LLMs to guide the decoding when generating codes, in which IDE-assistance is the intelligent features provided by Language Server. Compared to this work, we combine static analysis with RAG to provide more contexts instead of guiding the decoding process.
@@ VIII Conclusions
We propose Synbciatr, an LLM-powered approach to automatically repair obsolete test cases caused by syntactic changes of the focal methods. The key idea of Synbciatr is to combine the static analysis and neural rerankers to precisely construct test-repair-oriented contexts from the updated repository. Experimental results show that Synbciatr’s effectiveness outperforms baseline approaches on both the textual- and intent-matching metrics. Besides, with the augmentation of TROCtx constructed by Synbciatr, hallucinations are reduced by 57.1%. The overall results demonstrate that adopting static analysis techniques to improve the capability of LLM yields excellent performance and could be extended to other code-related tasks.

### v2

@@ V-A2 Human Evaluation on Effectiveness
Since textual metrics focus on measuring the similarity of tokens and AST structure for the given codes, they can not well represent the correctness of repair. To bridge this gap, in this part, we replace the obsolete test cases with the generated repaired ones in the repository and manually check whether the replaced test is correct or not, where a correct repair should pass the compilation and keep the intent of the test unchanged.
@@
Synter | 96.2% | 75.0% (102) | 90.4% (123) |
The column CPR in Tab. II represents the success rate of compilation after repairing the original test case in the repository. As a result, Synter fixes the most compilation errors caused by SynBCs. In terms of the two metrics of repairability, Synter both outperforms the baselines. As shown in the column Intent Match, we can observe that Synter correctly repair 75.0% (102/136) and 90.4% (123/136) test cases in alignment with the intent of the ground truth and the original test respectively.
As for specifically, Synter achieves improvements of 248.6% and 9.8% when compared to Ceprot and Naivellm respectively.
@@ V-B RQ2: Effectiveness of TROCtx
As introduced in Section I, Synter constructs TROCtxs to enhance LLMs, which serves as contextual information to reduce hallucinations during the generation of LLM (LLM below all refers to GPT-4). In this research question, we focus on illustrating the effectiveness of TROCtxs. Therefore, we compare our proposed Synter with Naivellm based on the remaining hallucinations.
@@
Common hallucinations. Common hallucinations are caused by the lack of direct contextual information from the change of focal method signature. For the case in Fig. 1, a new class (MountPOptions) is set as the third parameter type. Therefore, the LLM falls into common hallucination without the required class contexts of MountPOptions.
Outdated hallucinations. Outdated hallucinations are caused by the lack of implicit contextual information beyond the focal method signature. For the example in Fig. 5, the method findUnknown used in the test is refactored to be accessed from an instance of class BitstreamFormat. Without this specific knowledge, the test generated by LLM still uses the method in an outdated way, which causes a hallucination.
To assess the effectiveness of TROCtxs constructed by Synter, we collect all the hallucinations for Naivellm and Synter. As shown in Tab. III, hallucinations occur in both approaches. For the 6 common hallucinations that Naivellm fails, Synter fixes all of them and correctly generates repaired tests with related contexts, indicating that Synter has the general capability to precisely and effectively collect required contexts from the syntactic change of the focal method. Besides, Synter also reduces outdated hallucinations by 25%, in which two cases are fixed with the constructed TROCtxs (UsageCtxs and EnvCtxs specifically).
@@ V-C RQ3: Failure Analysis
Synter is designed to repair the original test case with its intent unchanged. Considering that the ground truth may contain semantic changes such as adding new assertions and altering input values, we further investigate the 13 cases that fail in terms of according to Tab. II. Finally, we summarize three reasons for the failure.
Uses unimported classes. For four cases, LLM directly uses classes that are not imported in the file of the test method when generating codes. The key reason is that we focus on collecting contexts by analyzing the syntactic change of the focal method at the method level. Without analyzing the focal class and changes in the methods, the constructed contexts by Synter may be inadequate and result in failure. Fortunately, this type of failure is easy to fix.
Complex focal changes. For eight cases, LLM fails to correctly generate the expected tests with limited contexts as the changes of focal methods in class-level or repo-level upgrades are complex. Specifically, the complex focal change leads to outdated hallucinations and incorrect invocations. This type is challenging to resolve because of the difficulty of collecting implicit contexts and the limited capability of current LLMs.
Fails to construct TROCtxs. Specifically for one case, we find that the language server fails to provide intelligent features for repositories containing configuration errors, which results in the failure of Synter to construct TROCtxs. Overall, it is hard even for developers to manually repair tests in these repositories, in which they can not infer related contexts in IDEs either.
@@ V-D RQ4: Efficiency
As shown in RQ1, the performance of Ceprot lags behind other approaches because of the limited model backbone. Although Ceprot can repair tests fast, most of the generated codes are incomplete and contain syntax errors (52.2%). Therefore, we focus on comparing the efficiency of Naivellm and Synter in this research question.
Compared to Naivellm, Synter adds the steps to construct TROCtxs. When querying LLM to repair the test, the prompt of Synter is longer with the constructed TROCtxs. In this research question, we evaluate the time efficiency and the token count of Synter compared to Naivellm.
In particular, we divide the process of Synter into constructing TROCtxs and querying LLM for repair. As shown in Fig. 6, the time cost of Synter mainly depends on the process of constructing TROCtxs, while the response time of LLM API of the two approaches is close. We also notice that the language server may fail to synchronize old repositories due to connection timeout for missing dependencies, which results in the outliers with much longer time. In our design, Synter trades time for accuracy. Besides, the average time of Synter falls into 10s to 30s, which is acceptable for developers since test repair is not an interactive task.
In terms of the token cost of Synter, we observe that the tokens of the prompt roughly doubled in number after being augmented with TROCtxs from Fig. 7. This evidence also indicates that the TROCtxs constructed by Synter are precise considering the large token size of the whole repository.
VI Discussion
VI-A Limitations
The key limitations of Synter exist in its two main modules, the static collector and the neural reranker. First, Synter cannot construct related contexts if the language server fails to initialize due to configuration errors in the repository. Second, the approach’s effectiveness is influenced by its reranking strategy. Despite crafting tailored reranking queries for different contexts, we cannot always ensure the precision of reranking. Additionally, our approach focuses on method-level signature changes and does not improve the performance of test repair for class-level or complex implicit changes specifically.
VI-B Threats to Validity
External Validity. The main threats to external validity come from the evaluation dataset. We reuse the existing dataset, which may not be representative of all possible real-world syntactic changes, so we check and augment the dataset to collect diverse samples from related commits. Besides, it is inevitable to avoid data leakage from popular open-source projects in GitHub. These projects widely use standard design patterns [38], which can be learned by LLM during training. Although experimental results show that Naivellm correctly repairs some of the tests with data leakage, it does not affect the effectiveness evaluation of Synter. According to the analysis in RQ2, Synter outperforms Naivellm in cases where Naivellm fails for hallucinations.
@@ VII-A Production-test Co-evolution
Two types of obsolete test cases have attracted the attention of developers, one is the GUI-oriented event sequence test case, and the other is the code-oriented method test case. Nowadays, many researchers studied the automated repair of GUI test cases, especially on Android applications [40], [12]. However, for code-oriented method test cases, limited studies are focusing on repairing code-oriented method test cases automatically. One biggest difference between them is that the search space for the GUI-oriented test event is more obvious, which can be obtained by traversing the related GUI widget trees, while the code-oriented method test case has a larger search space in the whole repository.
For code-oriented method test case repairing, Hu et al. [13] proposed the first transformer-based approach to update obsolete tests with two stages, identifying and updating, which is the SOTA work in this area. However, it is based on pre-trained models with fewer parameters and lacks contextual information. Compared to this, our work uses larger language models with automatically constructed contexts from the repository.
@@ VII-B LLM-based Code Generation
Automated code generation with LLMs brings huge improvements in production efficiency. Repo-level code generation represents the task of generating codes based on a broader context of the repository [41]. It is challenging due to the lack of domain-specific knowledge, which results in hallucination [42], [43]. Several studies leverage Retrieval-Augmented Generation (RAG) to improve the performance of LLM in specific code tasks by providing similar codes or results into the query prompt [41], [26]. However, the contexts are retrieved based on simple metrics such as textual similarity without code semantics. Recently, some works have focused on improving the capability of LLM by static analysis [44], [20]. Specifically, monitor-guided decoding (MGD) [20] is a novel approach to bring IDE-assistance from developers to LLMs to guide the decoding when generating codes, in which IDE-assistance is providing intelligent features by the language server. Compared to this work, we augment LLMs by combining static analysis with RAG to provide more contexts instead of guiding the decoding process.
@@ VIII Conclusions
We propose Synter, an LLM-powered approach to automatically repair obsolete test cases caused by syntactic changes of the focal methods. The key idea of Synter is to combine static analysis and neural rerankers to precisely construct test-repair-oriented contexts from the updated repository, which augments the capability of LLM. Experimental results show that Synter’s effectiveness outperforms baseline approaches on both the textual- and intent-matching metrics. Besides, with the augmentation of TROCtx constructed by Synter, hallucinations are reduced by 57.1%. The overall results also demonstrate that adopting static analysis techniques to improve the capability of LLM yields excellent performance and could be extended to other code-related tasks.
Acknowledgements
Thanks to Dr. Xutong Ma for the initial discussions on this work and to all the anonymous reviewers for their helpful comments and suggestions. This work is supported by the National Natural Science Foundation of China (NSFC) under grant numbers 62132020 and 62102405, and Major Project of ISCAS (ISCAS-ZD-202302).